### v1

# Vulnerabilities Unveiled: Adversarially Attacking a Multimodal Vision Langauge Model for Pathology Imaging

Abstract
In the dynamic landscape of medical artificial intelligence, this study meticulously explores the vulnerabilities of the Pathology Language-Image Pretraining (PLIP) model, a Vision Language Foundation model, under targeted adversarial conditions. Leveraging the Kather Colon dataset with 7,180 H&E images across nine tissue types, our investigation employs Projected Gradient Descent (PGD) adversarial attacks to intentionally induce misclassifications. The outcomes reveal a 100% success rate in manipulating PLIP’s predictions, underscoring its susceptibility to adversarial perturbations. The qualitative analysis of adversarial examples delves into the interpretability challenges, shedding light on nuanced changes in predictions induced by adversarial manipulations. These findings contribute crucial insights into the interpretability, domain adaptation, and trustworthiness of Vision Language Models in medical imaging. The study emphasizes the pressing need for robust defenses to ensure the reliability of AI models.
[…]
Index Terms—  Adversarial Attacks, Histopathology Data, Vision Language Foundation Models, Pathology, AI, Robustness, Trustworthiness, Medical Image Analysis
1 Introduction
Fig. 1: Attack Overview
The incorporation of artificial intelligence (AI) into the field of medical imaging and pathology has experienced notable advancements, leading to substantial progress in the areas of diagnoses and analysis  [1] [2]. Furthermore, the utilization of artificial intelligence (AI) in the field of pathology goes beyond traditional imaging methods and encompasses the intricacies associated with Hematoxylin and Eosin (H&E) staining  [3]. This integration allows for a thorough comprehension of tissue structures and the morphology of cells.
The Vision Language Foundation concept, represented by the Pathology Language-Image Pretraining (PLIP) approach, is positioned at the forefront of this paradigm change. PLIP is an AI framework that operates in several dimensions, effectively managing the complex interplay between visual and textual data  [4]. It leverages the collective knowledge found on platforms such as medical Twitter to overcome the limitations posed by a scarcity of annotated medical images. This significant advancement not only tackles the issue of limited data availability but also enables pioneering research in the field of pathology analysis. These models have exhibited exceptional performance in zero-shot classification, representing the cutting-edge capabilities in this domain.
Fig. 2: Original H&E + Pertubation = Adversarial H&E
Nevertheless, with the increasing integration of AI models like PLIP into pathology investigation, the potential threat of adversarial attacks becomes a significant concern. The reliability of AI systems in medical imaging can be compromised by many techniques, such as Fast Gradient Sign Method (FGSM) attack  [5], Carlini & Wagner  [6], and projected gradient descent (PGD) attack  [7]. This technique has the potential to pose significant challenges to the integrity and accuracy of AI systems in this domain.
In the field of pathology, the utmost importance is placed on the factors of interpretability and trustworthiness. The lack of transparency and comprehensibility in the decision-making process of numerous machine learning models gives rise to issues regarding their black-box character and the capacity to understand the underlying reasoning behind their forecasts.
In this domain, the ability to make informed decisions relies heavily on the comprehensibility of the outputs generated by AI systems. This research study examines the complex domain of artificial intelligence (AI) in the field of medical imaging, specifically focusing on the details and complexities of the PLIP model as a foundational framework for vision-based language processing. By conducting a thorough investigation of adversarial attacks and emphasizing the need of interpretability and trustworthiness, our objective is to provide significant contributions to the ongoing academic discussion regarding the crucial impact of AI on the development of pathology analysis in the future.
2 methods
2.1 Threat Model
When considering the challenges posed by opposing forces, our examination of the threat model centers on the reliability of Vision Language Models (VLMs), specifically the Pathology Language-Image Pretraining (PLIP). Consider hypothetical situations in which diagnostic results are altered with the intention of obtaining financial benefits, hence introducing unethical components and the possibility of exploitation. In addition to financial motivations, we explore the deliberate manipulation of diagnostic outcomes by opponents, motivated by monetary incentives or a quest for market dominance within the healthcare industry.
2.2 Vision Language Models - PLIP
Our current investigation focuses on the methodological aspects of the Pathology Language-Image Pretraining (PLIP) model, including examining its architectural intricacies. PLIP has been chosen based on its exceptional ability to effectively include visual and textual data, making it an advanced Vision Language Model (VLM) that holds a distinctive position in the field of pathology image analysis. It achieves this by utilizing the collective wisdom shared on platforms such as medical Twitter to carefully select and organize the OpenPath dataset. This dataset is a substantial resource consisting of 208,414 pathology images accompanied by descriptions in natural language. This model has been rigorously evaluated on popular datasets such as KatherColon  [8], PanNuke  [9], DigestPath  [10] datasets for zero-shot image classification.
Our objective in choosing PLIP is to provide detailed perspectives on the reliability and resilience of Vision Language Models, such as PLIP, in the crucial field of pathology analysis. This exploration functions as a crucial link between state-of-the-art technology and its tangible applications in real-life situations.
2.3 Dataset
Fig. 3: PieChart of Kather Colon Dataset
Our study relies on the Kather Colon dataset, a carefully selected 7,180 image patches from 50 colorectal adenocarcinoma patients. The dataset in question functions as a critical tool for assessing the performance and vulnerability of PLIP in the face of adversarial attacks. The dataset used in this study consists of photographs with dimensions of 224x224 pixels and a resolution of 0.5 megapixels per pixel (MPP). It serves not only as a testing platform but also as a validation set with clinical relevance, hence assuring the practicality and application of our research findings in real-world scenarios. Fig.3. is a piechart representing 9 tissue types with the number of images in each tissue type.
2.4 Adversarial Attack - Projected Gradient Descent
We chose to perform adversarial attack on Pathology Language Image Pretraining (PLIP) by performing The Projected Gradient Descent (PGD) method. The choice of PGD is supported by its efficacy as an iterative optimization method, enabling the creation of perturbations that are both effective and subtle. This particular attribute is in perfect harmony with our objective of methodically examining and comprehending the vulnerabilities of PLIP under controlled hostile circumstances. This enables a thorough evaluation of the resilience of PLIP against sophisticated attacks.
As shown in Fig. 1 we take the original H&E image and perform Projected Gradient Descent (PGD) attack by iterating through several steps to achieve optimal pertrubation. We perform this attack based on two main objectives, which are targeted misclassification and a high Structural Similarity Index Measure (SSIM) score.
3 Experiements and Results
3.1 Evaluation
In order to provide a clearer understanding of the influence of adversarial attacks on the predictions made by PLIP, we utilized heat maps as a visual tool to depict the model’s label predictions both prior to and subsequent to the attacks. See Fig. 4 A continuous pattern of accurate label predictions is observed in the top heat map obtained from the original photos. However, the observed pattern is significantly disturbed in the associated adversarial heat maps on the botton of Fig. 4. This investigation highlights the difficulty presented by adversarial manipulations, placing emphasis on the significance of interpretability in medical artificial intelligence (AI) models.
To strengthen our assessment, we calculated the Structural Similarity Index (SSIM) ratings for both the unaltered and manipulated images. Notably, it was noted that the Structural Similarity Index (SSIM) values regularly above 90%, suggesting a significant level of resemblance between the unaltered and modified photos. The targeted PGD attacks aimed at inducing misclassifications within PLIP’s predictions for specific tissue types were remarkably successful, yielding a 100% attack success rate. From Fig. 5, we observe that we have reached 100% attack sccess rate for all the tissue types after 10 steps to achieve optimal pertubation. This shows how vulnerable the vision language models are.
Fig. 4: Heatmaps showing distribution of Adversarial Attacks
3.2 Targeted Misclassification
In the context of our targeted PGD adversarial attacks on the Pathology Language-Image Pretraining (PLIP) model using the Kather Colon dataset, our objective was to intentionally induce misclassifications for specific tissue types.
These targeted misclassifications were chosen to assess the model’s susceptibility to adversarial manipulations across a spectrum of tissue types, mirroring potential real-world scenarios where misdiagnoses could have critical consequences. The successful implementation of these targeted misclassifications, which can be observed from Fig. 2, further underscores the nuanced vulnerabilities of PLIP to adversarial attacks and raises questions about the model’s trustworthiness in medical imaging applications.
Fig. 5: Attack Success Rate per step
4 Conclusion and Future Directions
In the pursuit of advancing the interpretability, domain adaptation, and trustworthiness of medical artificial intelligence, our investigation into the vulnerabilities of the Pathology Language-Image Pretraining (PLIP) model through targeted PGD adversarial attacks reveals critical insights. The 100% success rate in inducing intentional misclassifications underscores PLIP’s susceptibility to adversarial manipulations, prompting a fundamental reassessment of its trustworthiness in medical imaging applications. The qualitative analysis of adversarial examples illuminates the intricate changes in PLIP’s predictions, emphasizing the need for interpretability-aware defenses to fortify the model against unforeseen challenges. These findings contribute to the broader discourse on the robustness of Vision Language Models in pathology analysis, guiding the development of AI models that not only exhibit high performance but also maintain reliability in the face of adversarial complexities.
5 Acknowledgments
This work was supported by a University of Texas System Rising STARs Award (J.M.L) and the CPRIT First Time Faculty Award (J.M.L)
6 Compliance with ethical standards
This research study was conducted retrospectively using human subject data made available in open access. Ethical approval was not required as confirmed by the license attached with the open access data.